OpenStudy (anonymous):

Simplify: (3a3b4)2(b3)3.

OpenStudy (callisto):

Is this your question? \[(3a^3b^4)^2(b^3)^3\]

OpenStudy (anonymous):

Yes.

OpenStudy (callisto):

\[(3a^3b^4)^2(b^3)^3\]\[=3^2 a^{3\times 2}b^{4\times 2 + 3 \times 3} = ...?\]

OpenStudy (anonymous):

?

OpenStudy (callisto):

Can you simplify it??

OpenStudy (anonymous):

Yes Thank you :D

OpenStudy (callisto):

welcome.. :)